OpenStudy (sphott51):

3 + sq rt of z-11 = - sq rt of z+4

OpenStudy (leenathan):

http://www.mathworks.com/help/symbolic/mupad_ref/sqrt.html?requestedDomain=www.mathworks.com o direct answers (this should help)

OpenStudy (sphott51):

the negative is on the outside of the sq rt so i'm not entirely sure what to do with it.

jhonyy9 (jhonyy9):

so there are 3 +sqrt(z-11) = -sqrt(z+4) yes ?

OpenStudy (sphott51):

yes

jhonyy9 (jhonyy9):

ok so like a first step square both sides so what will get ?

OpenStudy (sphott51):

3 +z-11 = (- ? ) z+4

jhonyy9 (jhonyy9):

no than square both sides will be (3 + sqrt(z-11) )^2 = (- sqrt(z+4) )^2

OpenStudy (sphott51):

oh so no negative?

jhonyy9 (jhonyy9):

on the lrft part use formula of (a+b)^2 = ?

jhonyy9 (jhonyy9):

on the left part

OpenStudy (sphott51):

9 + z-11 = z+4

jhonyy9 (jhonyy9):

on the left part not is right (a+b)^2 = a^2 +2ab +b^2

OpenStudy (sphott51):

Erm I basically just crossed out the square and square root 3^2 is 9

jhonyy9 (jhonyy9):

but so not is correct bc, you need square the all right side with these two terms and so will need using this formula of (a+b)^2

jhonyy9 (jhonyy9):

sory left side not right side

OpenStudy (sphott51):

Okay, sorry i'm not getting it?

jhonyy9 (jhonyy9):

but i ve wrote there above all what you need to use

OpenStudy (sphott51):

So I literally have to do the square root part twice?

jhonyy9 (jhonyy9):

on the left side use formula of (a+b)^2 = a^2 +2ab +b^2 with a=3 and b=sqrt(z-11)

OpenStudy (sphott51):

\[\sqrt{z-11}^2\]

OpenStudy (sphott51):

I don't know how to do this